Assume you are the manager of a local hospital. You begin the year in January with 150 boxes of inventory, each at a cost of $1,000. Your organization had the following inventory changes during the year: You purchased 110 additional boxes in March at a cost of $1,200 each, purchased an additional 230 boxes in August at a cost of $1,400 each, and used 400 boxes during the year. Calculate the value of the ending inventory and the value of the inventory used for the year using both the FIFO and LIFO methods


1
Expert's answer
2021-03-29T12:00:41-0400

The value of the ending inventory and the value of the inventory used for the year using the FIFO method are:

Value of the inventory used = 150×1,000 + 110×1,200 + 140×1,400 = $478,000.

Value of the ending inventory = 90×1,400 = $126,000.

The value of the ending inventory and the value of the inventory used for the year using the LIFO method are:

Value of the inventory used = 230×1,400 + 110×1,200 + 60×1,000 = $514,000.

Value of the ending inventory = 90×1,000 = $90,000.